solve by using the addition method
x-y=100
0.20x+0.06y=150

Answer by Abbey(339)

I think you mean solve by using the elimination method...
x -y = 100
.2x +.6y = 150
Multiply the second equation by -5
x- y = 100
-x+-3y =-750
Add the two equations to get
-4y=-650
Divide both sides by -4
y=162.5
Substitute 325 into first equation to get
x-162.5=100
Add 162.5 to both sides
x=262.5
check your answer with the second equation:
.2(262.5)+.6(162.5)=150
52.5+97.5=150 (True statement)
The solution is (262.5,162.5)